OpenStudy (anonymous):

(I DONT EVEN KNOW WHAT THIS IS PLEASE HELP ME!!) Evaluate. A.)9!/6!3! B.)1 C.)84 168

OpenStudy (uri):

Probability?

OpenStudy (anonymous):

evaluate - 9!/6!3! A. 1 B.84 C.168

OpenStudy (anonymous):

I guess?

OpenStudy (whpalmer4):

\[n! = n*(n-1)*(n-2)...(2)(1)\] For example,\[5!= 5*4*3*2*1\]

OpenStudy (anonymous):

how? why...??

OpenStudy (anonymous):

so how the heck do I do this one??

OpenStudy (whpalmer4):

how? why..?? what? you're supposed to evaluate \[\frac{9!}{6!3!} = \frac{9*8*7*6*5*4*3*2*1}{(6*5*4*3*2*1)(3*2*1)} =\]

OpenStudy (whpalmer4):

some careful cancellation will be most helpful!

OpenStudy (uri):

9*8*7*6*5*4*3*2*1/6*5*4*3*2*1x 3*2*1=362880/720*6=84

OpenStudy (anonymous):

..... o.o....... uhm.....ima just give up and guess.... im sorry for wasting yalls time. Bye.

OpenStudy (whpalmer4):

Why are you giving up? This is easy! \[\frac{9!}{6!3!} = \frac{9*8*7*6*5*4*3*2*1}{(6*5*4*3*2*1)(3*2*1)} = \frac{9*8*7*\cancel{6*5*4*3*2*1}}{\cancel{(6*5*4*3*2*1)}(3*2*1)}\]\[ = \frac{9*8*7}{3*2*1} =\]

OpenStudy (anonymous):

for you. im 100% dyslexic in algebra and both of my algebra teachers have been the worst out of the entire school. ^^

OpenStudy (whpalmer4):

Just write down the numbers, like I said. Cross out pairs top and bottom. Multiply and divide what is left. Don't let the appearance scare you off.

OpenStudy (uri):

Isn't 84 the answer? @whpalmer4

OpenStudy (whpalmer4):

@uri yes, it is.

OpenStudy (uri):

Thaankgod,I know my statistics.Thanks to @terenzreignz ♥

OpenStudy (whpalmer4):

\[\frac{9*8*7}{3*2*1} = \frac{504}{6} = 84\]alternatively\[\frac{9*8*7}{3*2*1} = \frac{9}{3}*\frac{8}{2}*\frac{7}{1} = 3*4*7 = 84\]